OpenStudy (anonymous):

what is the measure of the complementary angle to 0.8 radians? Give an exact answer

OpenStudy (ybarrap):

Complementary angles sum to 90 degrees, right? So 0.8 radians is how many degrees? Or keeping this in radians, 0.8 plus what is \(\cfrac{\pi}{2}\), which is 90 degrees in radians? $$ \cfrac{8}{10}+x=\cfrac{\pi}{2} $$ Solve for x. That's it!

OpenStudy (anonymous):

x=-4/5+pi/2?

OpenStudy (ybarrap):

That's right! $$ x=-\cfrac{4}{5}+\cfrac{\pi}{2}=-\cfrac{8}{10}+\cfrac{5\pi}{10}=\cfrac{5\pi-8}{10} $$ Which is an exact answer.